# GovBench: Benchmarking LLM Agents for Real-World Data Governance Workflows

### v1

@@ Abstract
Data governance ensures data quality, security, and compliance through policies and standards—a critical foundation for scaling modern AI development. Recently, Large Language Models (LLMs) have emerged as a promising solution for automating data governance by translating user intent into executable transformation code. However, existing benchmarks for automated data science often emphasize snippet-level coding or high-level analytics, failing to capture the unique challenge of data governance: ensuring the correctness and quality of the data itself. To bridge this gap, we introduce GovBench, a benchmark featuring 150 diverse tasks grounded in real-world scenarios, built on data from actual cases. GovBench employs a novel "reversed-objective" methodology to synthesize realistic noise and utilizes rigorous metrics to assess end-to-end pipeline reliability. Our analysis on GovBench reveals that current models struggle with complex, multi-step workflows and lack robust error-correction mechanisms. Consequently, we propose DataGovAgent, a framework utilizing a Planner-Executor-Evaluator architecture that integrates constraint-based planning, retrieval-augmented generation, and sandboxed feedback-driven debugging. Experimental results show that DataGovAgent significantly boosts the Average Task Score (ATS) on complex tasks from 39.7 to 54.9 and reduces debugging iterations by over 77.9% compared to general-purpose baselines.
GovBench: Benchmarking LLM Agents for Real-World Data Governance Workflows
@@ 1 Introduction
To address this limitation, we introduce GovBench, a hierarchically designed benchmark for natural-language-driven data governance. It contains 150 real-world tasks (100 operator-level; 50 DAG-level) covering six scenarios: Filtering, Refinement, Imputation, Deduplication & Consistency, Data Integration, and Classification & Labeling. GovBench’s key innovations include: i) an exceptionally comprehensive range of tasks, covering both atomic and multi-step processes, tasks on structured and unstructured data, as well as six distinct tasks grounded in real-world production environments. ii) a novel “reversed‑objective" methodology—that inverts the original task goal to programmatically generate task‑specific noise—to synthesize realistic and measurable noise; and iii) auto-generated, task-specific evaluation scripts that provide accurate normalized scores and standardized metrics to compare ground truth dataset with processed dataset—Code Runnable Rate (CRR), Task Success Rate (TSR), and Average Task Score (ATS)—ensuring a principled and reproducible assessment.
However, a robust benchmark is only half of the solution. When evaluated on GovBench, we find that even SOTA LLMs (openai2025gpt5; deepseekv3; openai2024gpt4o) baselines and general-purpose agent frameworks (qian-etal-2024-chatdev; li2023camel) exhibit a significant performance gap. They struggle to decompose complex instructions, generate logically correct multi-step pipelines, and debug errors, resulting in low task success rates. This indicates that, for data governance workflows, their architectures are not yet equipped with robust planning, strong grounding in established practices, or systematic debugging mechanisms.
@@
On GovBench-150, DataGovAgent significantly outperforms strong baselines. With GPT-5, it raises the Task Success Rate (TSR)—the core metric representing the proportion of tasks where business objectives are fully achieved—to 64% on operator-level tasks (+15 pp) and 60% on DAG-level tasks (+14 pp). Furthermore, compared to the strongest baseline ChatDev, DataGovAgent demonstrates superior efficiency by drastically reducing Average Debug Iterations (ADI) from 14.89 to 3.29.
@@
  • We introduce GovBench, the first hierarchical-designed benchmark for data governance automation, which features 150 realistic tasks based on real-world sources, noisy data synthesis and a rigorous, multi-metric evaluation protocol to address the critical gap in assessing end-to-end pipeline correctness.
@@ 2 Related Works
Current research emphasizes end-to-end workflow automation with minimal human intervention (sun2024survey). TheAgentCompany (xu2025theagentcompany) benchmarks LLM agents on consequential real-world tasks, while comprehensive surveys (baek2025automation; wang2025survey) highlight the transition from automation to autonomy in scientific discovery. These systems integrate planning, reasoning, reflection, and multi-agent collaboration capabilities. However, specialized data governance benchmarks remain limited. This gap highlights the necessity for benchmarks like our proposed GovBench.
@@
3 GovBench: A New Benchmark for Data Governance Automation
Figure 1: Distribution of the 100 operator-level tasks in GovBench across six governance categories: Filtering (22), Refinement (18), Imputation (18), Deduplication & Consistency (15), Integration (18), and Classification & Labeling (9). The split balances coverage of common governance operations while slightly emphasizing filtering and transformation tasks.
GovBench is a hierarchically designed data science benchmark dedicated to evaluating models’ capabilities in performing data governance tasks. It comprises 150 real-world data governance problems, including 100 operator-level tasks and 50 DAG-level tasks. GovBench comprehensively covers common scenarios encountered in real-life data governance workflows, including Filtering, Refinement, Imputation, Deduplication & Consistency, Data Integration, and Classification & Labeling. For each carefully curated NL task description, we proposed a reliable pipeline to synthesize ground-truth data and noisy data, accompanied by customized evaluation scripts to ensure precise and normalized scoring.
Figure 2: Illustration of the semi-automated pipeline designed for building GovBench, including real-world source data curation, hierarchical task objective design, noisy data synthesis, task-specific evaluation and DAG-level task construction.
@@ 3.2 Hierarchical Task Objective Design
GovBench comprises 100 Operator-level tasks and 50 DAG-level tasks. For Operator-level tasks, we designed six scenarios commonly encountered in real-world data governance, including Filtering, Refinement, Imputation, Deduplication & Consistency, Data Integration, and Classification & Labeling. We recruited five postgraduate students majoring in Computer Science, each with significant experience in data analysis and data science, to annotate Operator-level task descriptions. We provided a clear definition of the task description schema in advance and conducted a thorough manual validation to ensure both accuracy and clarity. As a result, 32% of the tasks were removed due to lack of feasibility, for instance, tasks with unclear objectives or ambiguous data requirements, and a final set of 100 task descriptions was retained. The distribution of tasks in these scenarios is illustrated in Figure 1.
@@ 3.4 Task-Specific Evaluation
Unlike existing benchmarks, which typically use numbers, strings, or similar forms as ground truth, GovBench uses original data sheets as its ground truth. Given the diverse nature of task categories and the varying characteristics of the source data, customized evaluation is essential for each task. Specifically, we designed a prompt template (See Prompt 4 to generate the task-specific evaluation scripts. Each task’s script compares the original data with the processed data and outputs a quantitative score between 0 and 1, reflecting the model’s effectiveness in completing the task. Evaluation logic is adjusted based on the specific nature of the task to ensure a precise assessment; a detailed breakdown for each Operator-level task category is provided in Table 6.
@@ 5 Experimental Setup
We evaluate on GovBench-150, which comprises 150 real-world data governance tasks. Each task includes a natural language description, raw datasets, and an executable script for automated scoring. The tasks are categorized into Operator-level (single-step operations) and DAG-level (complex multi-step workflows). We adopt the metrics detailed in Table 8, primarily focusing on Task Success Rate (TSR) for quality and Average Debug Iterations (ADI) for efficiency.
@@ 6.1 Performance of Single-Model Baselines
Table 1: Performance of Open-Source Models on GovBench (Operator-Level)
@@
Table 2: Performance of Closed-Source Models on GovBench (Operator-Level)
@@
Significant Performance Ceiling: Even the most powerful closed-source models, such as GPT-5 and Claude4-sonnet, fail to exceed a 50% TSR in a single-round code generation setting. This indicates that the tasks in GovBench are considerably challenging and difficult to solve perfectly with a single code generation attempt.
@@
Table 3: Performance of Open-Source Models on GovBench (DAG-Level)
@@
Table 4: Performance of Closed-Source Models on GovBench (DAG-Level)
@@ 6.2 Performance of Agent Framework Baselines
We evaluated the ChatDev and CAMEL frameworks on GovBench by pairing them with powerful GPT-4o and GPT-5 models in Figure 4.
Closing the Runnable–Correct Gap with Contracts and Feedback: On GovBench, DataGov‑Agent consistently turns runnability into business‑correct solutions more efficiently than generic agent frameworks. On DAG‑level tasks, although ChatDev+GPT‑5 attains the top TSR (64), DataGov‑Agent+GPT‑5 delivers higher average quality (ATS 54.91 vs. 39.67; Avg. Score 62.97 vs. 61.89), requires 4.5× fewer debug iterations (ADI 3.29 vs. 14.89). On operator‑level tasks, DataGov‑Agent+GPT‑5 leads in TSR/ATS/Avg. Score (64/55.47/69.15) and shows the strongest alignment between runnability and correctness (A=TSR/CRR=0.73 vs. 0.62 for ChatDev and 0.37 for CAMEL), indicating that contracts and meta‑cognitive feedback effectively convert CRR into TSR. More detailed analysis in Appendix A.5.
@@ 7 Conclusion
We present GovBench, the first benchmark designed to comprehensively stress-test large language model agents on real-world data governance tasks. GovBench offers two main contributions: it provides a two-tiered task suite that spans from atomic operators to multi-step DAG pipelines, and for each task, it incorporates unique evaluation logic and scoring metrics. Furthermore, our proposed DataGovAgent achieves SOTA performance on this new benchmark, significantly outperforming existing agent frameworks on complex governance pipelines.
@@ A.1 Benchmark Comparison Table
GovBench Hierarchical (Operator & DAG-level) 150 realistic tasks; reversed-objective noise; multi-metric scoring (ATS/TSR/CRR) End-to-end data-governance pipeline evaluation
@@ A.2 Tasks Eval
See evaluation Metrics for Operator-Level Task Categories in GovBench in Table 6.
@@ A.6 Ablation Study
To dissect the contribution of each component within the DataGovAgent framework, we conducted a series of ablation studies on the GovBench Operator-level tasks. We systematically disabled or replaced key modules—the Planner and the RAG mechanism to quantify their impact on overall performance. All experiments were run using GPT-5 as the base model. The results are summarized in Table 7.
Table 7: Ablation study of DataGovAgent on GovBench operator-level tasks. Numbers in brackets show the change ($\Delta$) w.r.t. the full model — red = decrease, green = increase.
@@ A.7 Metrics
Table 8: Evaluation Metrics for GovBench

### v2

# DataGovBench: Benchmarking LLM Agents for Real-World Data Governance Workflows

@@ Abstract
Data governance ensures data quality, security, and compliance through policies and standards—a critical foundation for scaling modern AI development. Recently, Large Language Models (LLMs) have emerged as a promising solution for automating data governance by translating user intent into executable transformation code. However, existing benchmarks for automated data science often emphasize snippet-level coding or high-level analytics, failing to capture the unique challenge of data governance: ensuring the correctness and quality of the data itself. To bridge this gap, we introduce DataGovbench, a benchmark featuring 150 diverse tasks grounded in real-world scenarios, built on data from actual cases. DataGovbench employs a novel "reversed-objective" methodology to synthesize realistic noise and utilizes rigorous metrics to assess end-to-end pipeline reliability. Our analysis on DataGovbench reveals that current models struggle with complex, multi-step workflows and lack robust error-correction mechanisms. Consequently, we propose DataGovAgent, a framework utilizing a Planner-Executor-Evaluator architecture that integrates constraint-based planning, retrieval-augmented generation, and sandboxed feedback-driven debugging. Experimental results show that DataGovAgent significantly boosts the Average Task Score (ATS) on complex tasks from 39.7 to 54.9 and reduces debugging iterations by over 77.9% compared to general-purpose baselines.
DataGovBench: Benchmarking LLM Agents for Real-World Data Governance Workflows
@@ 1 Introduction
To address this limitation, we introduce DataGovbench, a hierarchically designed benchmark for natural-language-driven data governance. It contains 150 real-world tasks (100 operator-level; 50 DAG-level) covering six scenarios: Filtering, Refinement, Imputation, Deduplication & Consistency, Data Integration, and Classification & Labeling. DataGovbench’s key innovations include: i) an exceptionally comprehensive range of tasks, covering both atomic and multi-step processes, tasks on structured and unstructured data, as well as six distinct tasks grounded in real-world production environments. ii) a novel “reversed‑objective" methodology—that inverts the original task goal to programmatically generate task‑specific noise—to synthesize realistic and measurable noise; and iii) auto-generated, task-specific evaluation scripts that provide accurate normalized scores and standardized metrics to compare ground truth dataset with processed dataset—Code Runnable Rate (CRR), Task Success Rate (TSR), and Average Task Score (ATS)—ensuring a principled and reproducible assessment.
However, a robust benchmark is only half of the solution. When evaluated on DataGovbench, we find that even SOTA LLMs (openai2025gpt5; deepseekv3; openai2024gpt4o) baselines and general-purpose agent frameworks (qian-etal-2024-chatdev; li2023camel) exhibit a significant performance gap. They struggle to decompose complex instructions, generate logically correct multi-step pipelines, and debug errors, resulting in low task success rates. This indicates that, for data governance workflows, their architectures are not yet equipped with robust planning, strong grounding in established practices, or systematic debugging mechanisms.
@@
On DataGovbench-150, DataGovAgent significantly outperforms strong baselines. With GPT-5, it raises the Task Success Rate (TSR)—the core metric representing the proportion of tasks where business objectives are fully achieved—to 64% on operator-level tasks (+15 pp) and 60% on DAG-level tasks (+14 pp). Furthermore, compared to the strongest baseline ChatDev, DataGovAgent demonstrates superior efficiency by drastically reducing Average Debug Iterations (ADI) from 14.89 to 3.29.
@@
  • We introduce DataGovbench, the first hierarchical-designed benchmark for data governance automation, which features 150 realistic tasks based on real-world sources, noisy data synthesis and a rigorous, multi-metric evaluation protocol to address the critical gap in assessing end-to-end pipeline correctness.
@@ 2 Related Works
Current research emphasizes end-to-end workflow automation with minimal human intervention (sun2024survey). TheAgentCompany (xu2025theagentcompany) benchmarks LLM agents on consequential real-world tasks, while comprehensive surveys (baek2025automation; wang2025survey) highlight the transition from automation to autonomy in scientific discovery. These systems integrate planning, reasoning, reflection, and multi-agent collaboration capabilities. However, specialized data governance benchmarks remain limited. This gap highlights the necessity for benchmarks like our proposed DataGovbench.
@@
3 DataGovbench: A New Benchmark for Data Governance Automation
Figure 1: Distribution of the 100 operator-level tasks in DataGovbench across six governance categories: Filtering (22), Refinement (18), Imputation (18), Deduplication & Consistency (15), Integration (18), and Classification & Labeling (9). The split balances coverage of common governance operations while slightly emphasizing filtering and transformation tasks.
DataGovbench is a hierarchically designed data science benchmark dedicated to evaluating models’ capabilities in performing data governance tasks. It comprises 150 real-world data governance problems, including 100 operator-level tasks and 50 DAG-level tasks. DataGovbench comprehensively covers common scenarios encountered in real-life data governance workflows, including Filtering, Refinement, Imputation, Deduplication & Consistency, Data Integration, and Classification & Labeling. For each carefully curated NL task description, we proposed a reliable pipeline to synthesize ground-truth data and noisy data, accompanied by customized evaluation scripts to ensure precise and normalized scoring.
Figure 2: Illustration of the semi-automated pipeline designed for building DataGovbench, including real-world source data curation, hierarchical task objective design, noisy data synthesis, task-specific evaluation and DAG-level task construction.
@@ 3.2 Hierarchical Task Objective Design
DataGovbench comprises 100 Operator-level tasks and 50 DAG-level tasks. For Operator-level tasks, we designed six scenarios commonly encountered in real-world data governance, including Filtering, Refinement, Imputation, Deduplication & Consistency, Data Integration, and Classification & Labeling. We recruited five postgraduate students majoring in Computer Science, each with significant experience in data analysis and data science, to annotate Operator-level task descriptions. We provided a clear definition of the task description schema in advance and conducted a thorough manual validation to ensure both accuracy and clarity. As a result, 32% of the tasks were removed due to lack of feasibility, for instance, tasks with unclear objectives or ambiguous data requirements, and a final set of 100 task descriptions was retained. The distribution of tasks in these scenarios is illustrated in Figure 1.
@@ 3.4 Task-Specific Evaluation
Unlike existing benchmarks, which typically use numbers, strings, or similar forms as ground truth, DataGovbench uses original data sheets as its ground truth. Given the diverse nature of task categories and the varying characteristics of the source data, customized evaluation is essential for each task. Specifically, we designed a prompt template (See Prompt 4 to generate the task-specific evaluation scripts. Each task’s script compares the original data with the processed data and outputs a quantitative score between 0 and 1, reflecting the model’s effectiveness in completing the task. Evaluation logic is adjusted based on the specific nature of the task to ensure a precise assessment; a detailed breakdown for each Operator-level task category is provided in Table 6.
@@ 5 Experimental Setup
We evaluate on DataGovbench-150, which comprises 150 real-world data governance tasks. Each task includes a natural language description, raw datasets, and an executable script for automated scoring. The tasks are categorized into Operator-level (single-step operations) and DAG-level (complex multi-step workflows). We adopt the metrics detailed in Table 8, primarily focusing on Task Success Rate (TSR) for quality and Average Debug Iterations (ADI) for efficiency.
@@ 6.1 Performance of Single-Model Baselines
Table 1: Performance of Open-Source Models on DataGovbench (Operator-Level)
@@
Table 2: Performance of Closed-Source Models on DataGovbench (Operator-Level)
@@
Significant Performance Ceiling: Even the most powerful closed-source models, such as GPT-5 and Claude4-sonnet, fail to exceed a 50% TSR in a single-round code generation setting. This indicates that the tasks in DataGovbench are considerably challenging and difficult to solve perfectly with a single code generation attempt.
@@
Table 3: Performance of Open-Source Models on DataGovbench (DAG-Level)
@@
Table 4: Performance of Closed-Source Models on DataGovbench (DAG-Level)
@@ 6.2 Performance of Agent Framework Baselines
We evaluated the ChatDev and CAMEL frameworks on DataGovbench by pairing them with powerful GPT-4o and GPT-5 models in Figure 4.
Closing the Runnable–Correct Gap with Contracts and Feedback: On DataGovbench, DataGov‑Agent consistently turns runnability into business‑correct solutions more efficiently than generic agent frameworks. On DAG‑level tasks, although ChatDev+GPT‑5 attains the top TSR (64), DataGov‑Agent+GPT‑5 delivers higher average quality (ATS 54.91 vs. 39.67; Avg. Score 62.97 vs. 61.89), requires 4.5× fewer debug iterations (ADI 3.29 vs. 14.89). On operator‑level tasks, DataGov‑Agent+GPT‑5 leads in TSR/ATS/Avg. Score (64/55.47/69.15) and shows the strongest alignment between runnability and correctness (A=TSR/CRR=0.73 vs. 0.62 for ChatDev and 0.37 for CAMEL), indicating that contracts and meta‑cognitive feedback effectively convert CRR into TSR. More detailed analysis in Appendix A.5.
@@ 7 Conclusion
We present DataGovbench, the first benchmark designed to comprehensively stress-test large language model agents on real-world data governance tasks. DataGovbench offers two main contributions: it provides a two-tiered task suite that spans from atomic operators to multi-step DAG pipelines, and for each task, it incorporates unique evaluation logic and scoring metrics. Furthermore, our proposed DataGovAgent achieves SOTA performance on this new benchmark, significantly outperforming existing agent frameworks on complex governance pipelines.
@@ A.1 Benchmark Comparison Table
DataGovbench Hierarchical (Operator & DAG-level) 150 realistic tasks; reversed-objective noise; multi-metric scoring (ATS/TSR/CRR) End-to-end data-governance pipeline evaluation
@@ A.2 Tasks Eval
See evaluation Metrics for Operator-Level Task Categories in DataGovbench in Table 6.
@@ A.6 Ablation Study
To dissect the contribution of each component within the DataGovAgent framework, we conducted a series of ablation studies on the DataGovbench Operator-level tasks. We systematically disabled or replaced key modules—the Planner and the RAG mechanism to quantify their impact on overall performance. All experiments were run using GPT-5 as the base model. The results are summarized in Table 7.
Table 7: Ablation study of DataGovAgent on DataGovbench operator-level tasks. Numbers in brackets show the change ($\Delta$) w.r.t. the full model — red = decrease, green = increase.
@@ A.7 Metrics
Table 8: Evaluation Metrics for DataGovbench